OpenStudy (anonymous):

Solve for h: h + 5 = h – 3 A. h = 8 B. h = -8 C. h = 2 D. no solution

OpenStudy (anonymous):

@blanca1

OpenStudy (johnweldon1993):

\[\large h + 5 =h - 3\] Lets think about that... does any number + 8 ....= that same number....- 3? Seems impossible to me

OpenStudy (anonymous):

D?

OpenStudy (johnweldon1993):

Correct...because for example...say that 'h' = 2 If we have \[\large h + 5 = h - 3\] \[\large 2 + 5 = 2 - 3\] \[\large 7 = -1\] See, it is impossible...so D is correct

OpenStudy (anonymous):

What is the y-intercept of the line? y = -3x A. -3 B. 0 C. 1 D. 3

OpenStudy (johnweldon1993):

A slope intercept form looks like \[\large y = mx + b\] Where 'm' = slope and 'b' = y-intercept Your equation is \[\large y = -3x\] Notice how you have the slope "-3" but there is no + b on there...so your y-intercept is 0 Because this is essentially like writing \[\large y = -3x + 0\]

OpenStudy (anonymous):

so always remember y+mx+B ?